Question #170676

Calculate the divergence of the following vector functions:

Vb = xy X +2yzY +3zx Z



1
Expert's answer
2021-10-05T15:43:51-0400

"\text{div }\vec V=\frac{\partial V_x}{\partial x}+\frac{\partial V_y}{\partial y}+\frac{\partial V_z}{\partial z}=y+2z+3x."